Allison:

The minute hand on a clock is 9 centimeters long and travels through an arc of 252 degrees every 42 minutes. To the nearest tenth of a centimeter, how far does the minute hand travel during a 42-minute period?

Allison:

@dude

dude:

 Draw it :P r is the minute hand aka 9cm greek letter is the angle it went to get to 42 minutes aka 252 \(s=2\pi r (\dfrac{\theta}{360})\)

Allison:

r is 9

Allison:

Greek letter 252

dude:

Yes

XluisX:

Shouldn't there be a decimal?

Allison:

39.58

dude:

Yes

Allison:

One more omg